# Pretrained deep models outperform GBDTs in Learning-To-Rank under label scarcity

### v4

@@ 5 Conclusion
We study the learning-to-rank (LTR) problem with tabular data under an abundance of unlabeled data, but with scarcity of labeled data—a very common scenario in real-world practical LTR systems. Prior works on supervised learning with tabular data have shown that GBDTs outperform deep learning (DL) models, especially on datasets with outliers. In this paper, we find that in the label-scarce setting, one can use DL pretraining methods, which can exploit available unlabeled data, to obtain new state-of-the art performance. Through experiments with public LTR datasets and a real-world large-scale online shopping dataset, we showed that pretraining allows DL rankers to outperform GBDT rankers, especially on outlier data. Finally, we provided empirically justified guidelines for pretraining on LTR datasets, including a new and faster LTR-specific pretraining method; SimCLR-Rank. There are still several open questions around pretraining for tabular and LTR data. For example, it is still an open question whether we can achieve knowledge transfer across different tabular datasets, as it is commonly done in image and text domains. It will also be useful to investigate if there are other LTR-specific pretraining methods which can further improve the ranking performance.
Acknowledgements
@@ A.1 Detailed discussion of related work
Deep tabular models. Given the success of neural methods in many other domains, there have been many attempts to adapt deep models to the tabular domain. TANGOS introduced special tabular-specific regularization to try to improve deep models’ performance (Jeffares et al., 2023). FT-transformer and TabTransformer were introduced as transformer-based approaches to tabular data (Gorishniy et al., 2021; Huang et al., 2020). All these models have failed to consistently or convincingly outperform tree-based methods based on their own evaluations.
@@
Inspired by the success of pretraining and self-supervised learning in images and text, several works show how to apply SSL to unlabeled tabular data. One strategy is to corrupt tabular data and train a deep model to reconstruct it (Yoon et al., 2020; Majmundar et al., 2022; Ucar et al., 2021; Nam et al., 2023b; Lin et al., 2023; Syed & Mirza, 2023; Hajiramezanali et al., 2022). Another approach is to use contrastive losses, which have been highly successful in the image domain (Chen et al., 2020). These methods are also applicable to our case because tabular data, like image data, is often composed of fixed-dimensional real vectors (Verma et al., 2021; Bahri et al., 2021; Lee & Shin, 2022; Hager et al., 2023; Liu et al., 2023; Darabi et al., 2021). Rubachev et al. (2022) evaluate a variety of different pretraining methods for tabular learning across many different datasets, finding that there is not a clear state of the art.
@@ A.2.1 Dataset statistics
In table Table 5, we provide the details of our three public datasets and the proprietary real-world dataset. Note that for the public dataset experiments we sub-sample the pre-training dataset in different ways (as described in Sections 3, 4.1.1, 3 and 4.1.2) to create the training dataset and simulate scarcity of labeled samples. However, we use the same validation dataset without and with the labels for the self-supervised pretraining and supervised learning stages, respectively. Note that in all these datasets, the higher the label value the higher the relevance of the item.
@@ A.2.2 Outliers query group details
Here we detail how we select the outlier query groups for the Outlier-NDCG metrics (Section 2.1). For our proprietary online shopping dataset, there are known outlier features that cause low quality predictions. These outliers arise from noise introduced in various stages in the data pipeline. We have mitigation strategies in place against these feature outliers, but they are often imperfect. Because we already know some of the outliers features, we deem any query group with items with outlier features as an outlier. These outliers comprise roughly 10% of the test set, which is still a large number of query groups (much larger than the number of query groups in the outlier datasets of the public datasets).
For our public datasets, we systematically select outliers as follows: we generate a histogram with 100 bins for each feature across the validation dataset. For example, Istella (Lamkhede & Kofler, 2021) has 220 features, so we have 220 different histograms. For each histogram, we scan from left to right on the bins until we have encountered at least $G$ empty bins in a row, and if there is less than 1% of the validation set above this bin, then all the feature values above this bin are considered outliers. We also repeat this process right to left. Any test query group containing items with outlier feature values is labeled an outlier query group, and placed in the test outlier dataset. In Figure 5, we present an example histogram of a feature in Istella_S dataset.
@@ A.2.3 Model Settings and Training Details
Implicit binary user feedback setting. We reuse the methodology of Section 4.1.1, except we use a scoring head of one linear layer (as opposed to a three layer MLP) for pretrained models. We found that this improved stability and performance in the binary label setting.
@@
GBDT + pseudolabeling. We add pseudo-labeled GBDTs as a semi-supervised GBDT baseline. For GBDTs with pseudo-labeling (GBDT-pseudo), we propose the following algorithm as there are no references to our knowledge. First, we train the GBDT on the labeled train set, and then use this GBDT to pseudo-labels for the unlabeled train set. A GBDT ranker outputs real-valued scores. To convert real-valued scores into relevance scores, we take the resulting scores and evenly bin them into 31 buckets. The scores converted to 0 are the lowest bucket, the scores converted to 1 are the second lowest bucket, and so on. We choose 31 because that is the largest number of labels allowed for GBDT ranking in lightgbm (the reason is because the LambdaRank loss requires the calculation of $2^{\textrm{label}}$ which can get very large).
@@
Comparing finetuning strategies. When linear probing (LP), we freeze the pretrained model and update only a linear head on top of it for 200 epochs. In multilayer probing (MP), we freeze the pretrained model and update a 3-layer MLP head on top of it for 200 epochs. In full finetuning (FF), we use the finetuning strategy from Section 4.1.1. We use the validation set for early stopping.
@@
A.2.4 Details of Binary Label generation used in Section 4.1.2
@@ A.3.1 Runtime comparisons of pre-training strategies
In Table 6, we provide the runtime comparisons between SimSiam, SimCLR, and SimCLR-Rank (new method).
@@ A.3.2 Additional Results on linear probing vs finetuning
Here we provide more results on the question of linear probing vs full finetuning, as discussed in Section 4.3. In Tables 7 and 8 we compare different probing strategies for all datasets when using SimSiam and SimCLR-Rank, respectively, as the pretraining method.
Table 7: SimSiam under different finetuning strategies on NDCG/Outlier-NDCG, averaged over 3 trials (LP = Linear Probing, MP = Multilayer Probing, FF = Full Finetuning). We find that linear probing and MLP probing perform extremely poorly (except in Yahoo Set1, where MLP probing performs well).
@@
FF 0.3149 $\pm$ 0.0119 0.52 $\pm$ 0.0133 0.6348 $\pm$ 0.0164
Table 8: SimCLR-Rank under different finetuning strategies on NDCG/Outlier-NDCG, averaged over 3 trials (LP = Linear Probing, MP = Multilayer Probing, FF = Full Finetuning). Multilayer probing performs moderately worse than full finetuning, and linear probing performs much worse than all other strategies.
@@ A.3.3 Additional Results on SimCLR-Rank vs SimSiam
In this section we give additional t-SNE plots of the embeddings learned by SimCLR-Rank and SimSiam encoders, and the pre-final layer of the fully supervised ranker (trained with all labels) on MSLRWEB30K dataset (Figure 6), and Istella_S dataset (Figure 7). For the Yahoo Set1 dataset, SimSiam performs the best, but for the above two datasets SimCLR-Rank is the best.
@@ A.3.4 Additional Results: Simulated crowdsourced label setting
Here we provide the results on label scarcity for Outlier-NDCG in Figure 8. Result interpretation is given in Section 4.1.1.
@@ A.3.5 Additional Results for Simulated Implicit Binary user feedback setting
In this part of the appendix we provide additional results on binary label generation as detailed in Section 4.1.2. In Tables 9 and 10 we give the dataset statistics when using $\tau_{\text{target}}=4.25$ and $\tau_{\text{target}}=5.1$, respectively. For these two settings, the GBDT baseline does not contain semi-supervised runs as we found that semi-supervision tends to generally degrade GBDT performance; see Section A.4. The “No Pretrain” baseline uses the tabular ResNet architecture.
@@
In Tables 12 and 13 we give the comparison between pretrained and non-pretrained rankers following the methodology in Section 4.1.2 and find that GBDTs perform the best when binary labels are less sparse ($\tau_{\text{target}}=4.25$) and pretrained DL rankers perform the best when binary labels are sparser ($\tau_{\text{target}}=5.1$). This strongly suggests that pretrained DL models outperform GBDTs when the labels arise from binary user feedback that is moderately uncommon or rare (as in online shopping, recommendations, or search).
Table 12: We compare pretrained models to non-pretrained models in the binary label setting with $\tau_{\text{target}}=4.25$ (Section 4.1.2) on NDCG averaged over three trials. We follow the methodology in Section 4.1.2.
@@
Table 13: We compare pretrained models to non-pretrained models in the binary label setting with $\tau_{\text{target}}=5.1$ (Section 4.1.2) on NDCG averaged over three trials. We follow the methodology in Section 4.1.2. indicates metrics on which pretrained rankers outperform non-pretrained rankers significantly via a $p<0.05$ t-test.
@@ A.3.6 Ablation: Data Augmentation Techniques
In Tables 14 and 15, we show the performance of each augmentation choice on SimCLR-Rank and SimSiam, respectively. We use the methodology described in Section 4.1.1, and have 0.2% of training QGs labeled.
Table 14: We compare different augmentation strategies for SimCLR-Rank. The methodology used is the one in Section 4.1.1, with 0.2% of training QGs labeled.
@@
Table 15: We compare different augmentation strategies for SimSiam. The methodology used is the one in Section 4.1.1, with 0.2% of training QGs labeled.
@@ A.3.7 Combining SimCLR-Rank and SimSiam
Here we describe how we constructively combined the SimCLR-Rank and SimSiam to get the competitive results in Section 4.3. Following the setting of Section A.2.3, we train (1) a SimCLR-Rank model with gaussian augmentation and noise of scale 1.0, and (2) a SimSiam model with zeroing augmentation and corruption probability 0.1.
@@ A.4 Semi-supervised GBDT
Here we provide a more detailed comparison between Supervised GBDTs, GBDT + pseudolabeling, and GBDT + PCA + pseudolabeling in Figure 9 for the simulated crowdsourcing of labels setting and in Table 16 for the simulated implicit binary user feedback setting. We follow the experimental methodology in Section A.2.3. We find that for the most part, semi-supervised learning does not benefit GBDT rankers.

### v5

@@ 5 Conclusion
We study the learning-to-rank (LTR) problem with tabular data under a scarcity of labeled data—a common scenario in real-world practical LTR systems. Prior works on supervised learning with tabular data have shown that GBDTs outperform deep learning (DL) models, especially on datasets with outliers. In this paper, we find that in the label-scarce setting, DL pretraining methods can exploit available unlabeled data to obtain new state-of-the art performance. Through experiments with public LTR datasets and a real-world large-scale online shopping dataset, we show that pretraining allows DL rankers to outperform GBDT rankers, especially on outlier data. Finally, we provide guidelines for pretraining on LTR datasets. There remain several open questions around pretraining for tabular and LTR data. It is still unknown whether we can achieve knowledge transfer across different tabular datasets, as is commonly done in image and text domains. It will also be useful to investigate LTR-specific pretraining methods which can further improve the ranking performance.
Acknowledgments
@@ A.1 Detailed discussion of related work
Deep tabular models. Given the success of neural methods in many other domains, there have been many attempts to adapt deep models to the tabular domain.
Borisov et al. (2022) categorize existing techniques for using deep neural networks over tabular data into four types: (1) Encoding-based methods such as VIME (Yoon et al., 2020), SCARF (Bahri et al., 2021), IGTD (Zhu et al., 2021), and SuperTML (Sun et al., 2019); (2) Novel hybrid architectures such as DeepFM (Guo et al., 2017), xDeepFM (Lian et al., 2018), and many others Cheng et al. (2016); Frosst & Hinton (2017); Ke et al. (2018; 2019); Popov et al. (2019); Luo et al. (2020); Liu et al. (2020); Ivanov & Prokhorenkova (2021); Luo et al. (2021); (3) Transformer-based architectures including SAINT (Somepalli et al., 2021), TabNet (Arik & Pfister, 2021), FT-Transformer Gorishniy et al. (2021), TabTransformer (Huang et al., 2020), and ARM-Net Cai et al. (2021)); (4) Regularized DNNs (Shavitt & Segal, 2018; Kadra et al., 2021). For instance, TANGOS introduced special tabular-specific regularization to try to improve deep models’ performance (Jeffares et al., 2023). We also note that a more recent work, Tree-hybrid simple MLP (Yan et al., 2024a), proposes to combine MLPs and GBDTs together for the tabular prediction problem.
@@
Inspired by the success of pretraining and self-supervised learning in images and text, several works show how to apply SSL to unlabeled tabular data. One strategy is to corrupt tabular data and train a deep model to reconstruct it (Yoon et al., 2020; Majmundar et al., 2022; Ucar et al., 2021; Nam et al., 2023b; Lin et al., 2023; Syed & Mirza, 2023; Hajiramezanali et al., 2022). Another approach is to use contrastive losses, which have been highly successful in the image domain (Chen et al., 2020). These methods are also applicable to our case because tabular data, like image data, is often composed of fixed-dimensional real vectors (Verma et al., 2021; Bahri et al., 2021; Lee & Shin, 2022; Hager et al., 2023; Liu et al., 2023; Darabi et al., 2021). Concurrent work has also investigated training tabular-specific LLMs for tabular tasks (Yan et al., 2024b; Schuh et al., 2024). Concurrent to us, Holzmüller et al. (2024) also proposes pretraining (with a modified architecture) to improve tabular deep learning performance. Rubachev et al. (2022) evaluate a variety of different pretraining methods for tabular learning across many different datasets, finding that there is not a clear state of the art.
@@
Pretrained models in learning to rank. Concurrent to the submission/review of this paper, recent works have also validated the benefits of pretraining in LTR. S2phere (Li et al., 2023c) proposes an elaborate recipe combining pseudo-labeling, contrastive learning, and transformer-based masked autoencoding to pretrain an LTR model, and shows that it performs well in offline and online metrics. MPGraf (Li et al., 2023b) proposes a pretraining method for LTR that combines both the query-document regression formulation of LTR and the graph-based formulation of LTR into a single model. GS2P (Li et al., 2024) proposes to use an ensemble of differently-trained models to generate pseudo-labels for the final model to learn. Like us, GRACE (Xu et al., 2024) uses contrastive learning for pre-ranking (retrieval); unlike our method, they use this process to align item embeddings with pretrained embeddings during supervised training phase. COLTR (Li et al., 2023d) proposes to increase the dimensionality of the features with Fourier features and to alternative pseudo-labeling between a listwise and pointwise ranker to produce better ranking results. Li et al. (2023a) finetune pretrained language models for the document ranking task. Ramazanli et al. (2025) study self-supervised learning for recommendation systems in concurrent work.
@@ A.2.1 Dataset statistics
In table Table˜5, we provide the details of our three public datasets and the proprietary real-world dataset. Note that for the public dataset experiments we sub-sample the pre-training dataset in different ways (as described in Sections˜3, 4.1.1, 3 and 4.1.2) to create the training dataset and simulate scarcity of labeled samples. However, we use the same validation dataset without and with the labels for the self-supervised pretraining and supervised learning stages, respectively. Note that in all these datasets, the higher the label value the higher the relevance of the item.
@@ A.2.2 Outliers query group details
Here we detail how we select the outlier query groups for the Outlier-NDCG metrics (Section˜2.1). For our proprietary online shopping dataset, there are known outlier features that cause low quality predictions. These outliers arise from noise introduced in various stages in the data pipeline. We have mitigation strategies in place against these feature outliers, but they are often imperfect. Because we already know some of the outliers features, we deem any query group with items with outlier features as an outlier. These outliers comprise roughly 10% of the test set, which is still a large number of query groups (much larger than the number of query groups in the outlier datasets of the public datasets).
For our public datasets, we systematically select outliers as follows: we generate a histogram with 100 bins for each feature across the validation dataset. For example, Istella (Lamkhede & Kofler, 2021) has 220 features, so we have 220 different histograms. For each histogram, we scan from left to right on the bins until we have encountered at least $G$ empty bins in a row, and if there is less than 1% of the validation set above this bin, then all the feature values above this bin are considered outliers. We also repeat this process right to left. Any test query group containing items with outlier feature values is labeled an outlier query group, and placed in the test outlier dataset. In Figure˜5, we present an example histogram of a feature in Istella_S dataset.
@@ A.2.3 Model Settings and Training Details
We also evaluate SAINT (Somepalli et al., 2021), a transformer-based pretrainining method. For the architecture, we follow (Somepalli et al., 2021), using $L=4$ layers, dropout of 0.1, attention heads $h=4$, with a self-attention query/key/value dimension of 16 and a intersample attention query/key/value dimension of 64. For augmentation, we follow (Somepalli et al., 2021) and use a CutMix mask parameter of 0.3 and MixUp parameter of 0.2. Again following (Somepalli et al., 2021) we weight the denoising loss 10 times more than the contrastive loss. For all other minor architectural details, we follow the code of (Somepalli et al., 2021). We pretrain for 5 epochs with a batch size of roughly 100 items due to GPU memory contraints and also the time cost (one epoch of saint took 300x longer than an epoch of SimCLR-Rank in MSLRWEB30K, 6000x longer in Yahoo Set1, and 1000x longer in Istella). We use a learning rate of 5e-5.
@@
For SAINT we follow the finetuning process of (Somepalli et al., 2021) in using the output of the CLS embedding as the representation for an item, and tune the learning rate in {5e-5, 1e-5, 5e-6} as we found the performance to be sensitive to this setting.
Implicit binary user feedback setting. We reuse the methodology of Section˜4.1.1, except we use a scoring head of one linear layer (as opposed to a three layer MLP) for pretrained models. We found that this improved stability and performance in the binary label setting.
@@
GBDT + pseudolabeling. We add pseudo-labeled GBDTs as a semi-supervised GBDT baseline. For GBDTs with pseudo-labeling (GBDT-pseudo), we propose the following algorithm as there are no references to our knowledge. First, we train the GBDT on the labeled train set, and then use this GBDT to pseudo-labels for the unlabeled train set. A GBDT ranker outputs real-valued scores. To convert real-valued scores into relevance scores, we take the resulting scores and evenly bin them into 31 buckets. The scores converted to 0 are the lowest bucket, the scores converted to 1 are the second lowest bucket, and so on. We choose 31 because that is the largest number of labels allowed for GBDT ranking in lightgbm (the reason is because the LambdaRank loss requires the calculation of which can get very large).
@@
SimCLR-Rank + GBDT/SimSiam + GBDT. We also explore using pretrained neural embeddings as features in the GBDT. To do this, we used the pretrained SimCLR-Rank (or SimSiam) model to produce embeddings over the labeled data, concatenated the original features with the embeddings, and then trained a GBDT on this new dataset. We use the same tuning grid as for the original GBDT model.
Comparing finetuning strategies. When linear probing (LP), we freeze the pretrained model and update only a linear head on top of it for 200 epochs. In multilayer probing (MP), we freeze the pretrained model and update a 3-layer MLP head on top of it for 200 epochs. In full finetuning (FF), we use the finetuning strategy from Section˜4.1.1. We use the validation set for early stopping.
@@
A.2.4 Details of Binary Label generation used in Section˜4.1.2
@@ A.3.1 Runtime comparisons of pre-training strategies
In Table˜6, we provide the runtime comparisons between SimSiam, SimCLR, and SimCLR-Rank (new method).
@@ A.3.2 Additional Results on linear probing vs finetuning
Here we provide more results on the question of linear probing vs full finetuning, as discussed in Section˜4.3. In Tables˜7 and 8 we compare different probing strategies for all datasets when using SimSiam and SimCLR-Rank, respectively, as the pretraining method.
Table 7: SimSiam under different finetuning strategies on NDCG/Outlier-NDCG, averaged over 3 trials (LP = Linear Probing, MP = Multilayer Probing, $F$ = Full Finetuning). We find that linear probing and MLP probing perform extremely poorly (except in Yahoo Set1, where MLP probing performs well).
@@
$F$ 0.3149 $\pm$ 0.0119 0.52 $\pm$ 0.0133 0.6348 $\pm$ 0.0164
Table 8: SimCLR-Rank under different finetuning strategies on NDCG/Outlier-NDCG, averaged over 3 trials (LP = Linear Probing, MP = Multilayer Probing, $F$ = Full Finetuning). Multilayer probing performs moderately worse than full finetuning, and linear probing performs much worse than all other strategies.
@@ A.3.3 Additional Results on SimCLR-Rank vs SimSiam
In this section we give additional t-SNE plots of the embeddings learned by SimCLR-Rank and SimSiam encoders, and the pre-final layer of the fully supervised ranker (trained with all labels) on MSLRWEB30K dataset (Figure˜6), and Istella_S dataset (Figure˜7). For the Yahoo Set1 dataset, SimSiam performs the best, but for the above two datasets SimCLR-Rank is the best.
@@ A.3.4 Additional Results: Simulated crowdsourced label setting
Here we provide the results on label scarcity for Outlier-NDCG in Figure˜8. Result interpretation is given in Section˜4.1.1.
@@ A.3.5 Additional Results for Simulated Implicit Binary user feedback setting
In this part of the appendix we provide additional results on binary label generation as detailed in Section˜4.1.2. In Tables˜9 and 10 we give the dataset statistics when using $\tau_{\text{target}}=4.25$ and $\tau_{\text{target}}=5.1$, respectively. For these two settings, the GBDT baseline does not contain semi-supervised runs as we found that semi-supervision tends to generally degrade GBDT performance; see Section˜A.4. The “No Pretrain” baseline uses the tabular ResNet architecture.
@@
In Tables˜12 and 13 we give the comparison between pretrained and non-pretrained rankers following the methodology in Section˜4.1.2 and find that GBDTs perform the best when binary labels are less sparse ($\tau_{\text{target}}=4.25$) and pretrained DL rankers perform the best when binary labels are sparser ($\tau_{\text{target}}=5.1$). This strongly suggests that pretrained DL models outperform GBDTs when the labels arise from binary user feedback that is moderately uncommon or rare (as in online shopping, recommendations, or search).
Table 12: We compare pretrained models to non-pretrained models in the binary label setting with $\tau_{\text{target}}=4.25$ (Section˜4.1.2) on NDCG averaged over three trials. We follow the methodology in Section˜4.1.2.
@@
Table 13: We compare pretrained models to non-pretrained models in the binary label setting with $\tau_{\text{target}}=5.1$ (Section˜4.1.2) on NDCG averaged over three trials. We follow the methodology in Section˜4.1.2. indicates metrics on which pretrained rankers outperform non-pretrained rankers significantly via a $p<0.05$ t-test.
@@ A.3.6 Ablation: Data Augmentation Techniques
In Tables˜14 and 15, we show the performance of each augmentation choice on SimCLR-Rank and SimSiam, respectively. We use the methodology described in Section˜4.1.1, and have 0.2% of training QGs labeled.
Table 14: We compare different augmentation strategies for SimCLR-Rank. The methodology used is the one in Section˜4.1.1, with 0.2% of training QGs labeled.
@@
Table 15: We compare different augmentation strategies for SimSiam. The methodology used is the one in Section˜4.1.1, with 0.2% of training QGs labeled.
@@ A.3.7 Combining SimCLR-Rank and SimSiam
Here we describe how we constructively combined the SimCLR-Rank and SimSiam to get the competitive results in Section˜4.3. Following the setting of Section˜A.2.3, we train (1) a SimCLR-Rank model with gaussian augmentation and noise of scale 1.0, and (2) a SimSiam model with zeroing augmentation and corruption probability 0.1.
@@ A.4 Semi-supervised GBDT
Here we provide a more detailed comparison between Supervised GBDTs, GBDT + pseudolabeling, and GBDT + PCA + pseudolabeling in Figure˜9 for the simulated crowdsourcing of labels setting and in Table˜17 for the simulated implicit binary user feedback setting. We follow the experimental methodology in Section˜A.2.3. We find that for the most part, semi-supervised learning does not benefit GBDT rankers.
A.5 Quantitative measures of embedding quality
We provide a quantitative measure of embedding quality in this section. We attempt to verify that the focus on in-query group negatives helps improve embedding quality. To do this, we compare the embeddings produced by a SimCLR-Rank model and a SimCLR-Sample model (SimCLR-Sample is the same as SimCLR-Rank except the negatives are randomly taken from the batch rather than being in-query group), both under zeroes augmentation with $p=0.1$, and measure how spread out the embeddings are within a query group for each method. In Table˜16, we find that SimCLR-Rank embeddings are consistently more spread out within a query group than SimCLR-Sample, showing that in-query group focus leads to a measurable impact on the distinguish-ability of embeddings within a query group. This may help explain why SimCLR-Rank performs better than SimCLR-Sample.
Table 16: The ratio of the in-query group mean absolute deviation over the overall mean absolute deviation of embeddings from SimCLR-Rank and SimCLR-Sample. These embeddings were calculated over the validation set of each dataset. We find that SimCLR-Rank has a higher ratio, meaning the embeddings are more spread out within a query group.
Method MSLRWEB30K Yahoo Set1 Istella
SimCLR-Rank 0.9953 0.9904 0.9915
SimCLR-Sample 0.9945 0.9881 0.9797